# MirrorCalib: Utilizing Human Pose Information for Mirror-based Virtual Camera Calibration

### v2

@@ IV-C1 Initial Estimation
Figure 4: Illustration of hip constraints. The left hip, right hip, and mid-hip keypoints should be co-linear.
IV-C2 Refining 2D Joint Locations
@@
Let $\mathbf{u_{i}^{t}}$ and $\mathbf{{u_{i}}^{t}}^{\prime}$ denote the 2D coordinates of joint $i$ and reflected joint $i^{\prime}$ at time $t$, $\mathbf{\Theta}$ and $\mathbf{\Theta^{\prime}}$ denote the extrinsic parameters of camera $C$ and the virtual camera $C^{\prime}$ respectively where $\mathbf{\Theta}$ is known. $\mathbf{\Theta^{\prime}}$ is updated by the modified eight-point algorithm in each iteration, and $\mathbf{X}_{i}^{t}$ denote the 3D coordinates of joint $i$ at time $t$. $\mathbf{X}_{i}^{t}$ can be calculated through triangulation from $\mathbf{u_{i}^{t}}$, $\mathbf{{u_{i}}^{t}}^{\prime}$, $\mathbf{\Theta}$ and $\mathbf{\Theta^{\prime}}$. Our goal is to “denoise” $\mathbf{u}_{i}^{t}$ and ${\mathbf{u}_{i}}^{t^{\prime}}$, to refine the estimation of $\mathbf{\Theta^{\prime}}$. One body prior is that the length of each limb remains constant over the duration of the recording. Let $l_{k}^{t}$ be the length of the $k_{th}$ bone in the bone set $P$ = {Right Femur, Left Femur, Left Humerus, Right Humerus, Left Ulna, Right Ulna, Left Tibia, Right Tibia, Left Scapular, Right Scaptular, Left Hip, Right Hip} at time $t$ from estimated 2D joint positions. To account for the discrepancies in bone length, we introduce a penalty term that considers the ratio of bone length variation over time to the mean bone length:
@@
Let $\mathbf{v_{ml}}^{t}$ ($\mathbf{v_{mr}}^{t}$) be the vectors from the left (right) hip to the mid-hip (Figure 4). It is expected that the left hip, right hip, and mid-hip joints are co-linear. Thus, we define:
@@
Figure 5: An epipolar plane is created by connecting a point in three-dimensional space to the optical centers of the real and virtual cameras, and this plane intersects with the image plane to form the corresponding epipolar line.
IV-C3 Outlier Rejection
@@
After the optimization in Section IV-C2, “denoised” 2D joint locations in T frames are obtained. There are many more pairs than needed to solve for the fundamental matrix. To further improve the calibration accuracy, we reject outliers by computing the epipolar distances for eligible pairs for a given fundamental matrix $F$. Specifically, given a point in three-dimensional space, an epipolar plane is defined by the intersection of the line connecting the point to the optical centers of two cameras as shown in Figure 5. This epipolar plane intersects the image planes and forms the epipolar lines. Let $l(x_{i},C^{\prime})=\mathbf{F}x_{i}$ and $l(x_{i}^{\prime},C)=\mathbf{F}^{T}{x_{i}}^{\prime}$ be the epipolar lines generated by the 2D point $x_{i}$ on the virtual view $C^{\prime}$ and the matched 2D point ${x_{i}}^{\prime}$ on the real view $C$ respectively. Then we can compute the sum of the distance from ${x_{i}}^{\prime}$ to $l(x_{i},C^{\prime})$ and the distance from $x_{i}$ to $l({x_{i}}^{\prime},C)$ in each frame in a video sequence and denote this sum as $g(\mathbf{F})=distance(l(x_{i},C^{\prime}),x_{i}^{\prime})+distance(l(x_{i}^{\prime},C),x_{i})$.
@@
Next, the Random Sample Consensus (RANSAC) algorithm for outlier removal is applied. Beginning with a set of six pairs of 2D joint positions selected at random, an iterative process is used to calculate the fundamental matrix. This matrix is then used to categorize the pairs into two groups: inliers and outliers, based on their respective epipolar distances. The largest inlier set is used to determine the fundamental matrix as described in Section IV-A. The detailed steps are summarized in Algorithm 1.
Algorithm 1 RANSAC Algorithm for Estimating Fundamental Matrix
1:
2:Corresponding feature points $\{x_{i},x^{\prime}_{i}\}_{i=1}^{n}$
3:The maximum number of iterations $K$
4:inlier threshold $\tau$
5:The minimum number of inliers $\gamma$
6:
7:Fundamental matrix $\mathbf{F}$
8:$itr\leftarrow 0$, $best\_inliers\leftarrow\varnothing$
9:while $itr<K$ and $|best\_inliers|<\gamma$ do
10:     $S\leftarrow$ random sample of 6 pairs of features from $\{x_{i},x^{\prime}_{i}\}_{i=1}^{n}$
11:     $\mathbf{F}\leftarrow$ compute fundamental matrix using $S$
12:     $inliers\leftarrow\varnothing$
13:     for $i=1$ to $n$ do
14:         if $d_{i}(\textbf{F})<\tau$ then
15:              Add $i$ to $inliers$
16:         end if
17:     end for
18:     if $|inliers|>|best\_inliers|$ then
19:         $best\_inliers\leftarrow inliers$
20:     end if
21:     $itr\leftarrow itr+1$
22:end while
23:$\mathbf{F}\leftarrow$ re-estimate fundamental matrix using all points in $best\_inliers$
24:return $\mathbf{F}$
V Evaluation
@@
In this section, we evaluate the performance of MirrorCalib on both synthetic dataset and real datasets. We conduct experiments to demonstrate the influence of the number of frames and the errors in 2D human pose estimation. Since MirrorCalib is dependent on the intrinsic parameters of cameras, the impact of the accuracy of intrinsic parameters is also evaluated. Furthermore, we conduct a case study on 3D HPE to illustrate the practical application of MirrorCalib.
V-A Implementation Details
V-A1 Intrinsic Parameters
@@
MirrorCalib assumes that the intrinsic parameters that characterize the focal length and optical center are given as in previous work [11, 12, 13]. These intrinsic parameters can be obtained from the metadata of videos recorded by professional cameras. Alternatively, they can be estimated using vanishing points [23] under the assumption that the camera has zero skew, square pixels, and the principal point is located at the image center. In the case where no ground-truth intrinsic parameters are available, we introduce additional steps to estimate these parameters in the proposed pipeline (Figure 2) as follows:
  1. 1.
    Use the detected 2D keypoints and mirror edges to estimate the focal length following the method in [23]. Specifically, the pairs of keypoints in the real and mirrored view are used to compute one vanishing points, while the second vanishing point is determined as the intersection of mirror edges from the real view;
  2. 2.
    Incorporate the estimated focal length in the normalized mirror-constrained eight-point algorithm (Section IV-A), to derive the initial set of extrinsic parameters;
  3. 3.
    Refine the extrinsic parameters through optimization, integrating prior knowledge of the human body;
  4. 4.
    Employ the RANSAC algorithm to identify inliers for the re-estimation of extrinsic parameters;
  5. 5.
    Repeat (1) using only the selected inliers and mirror edges;
  6. 6.
    Compute the final set of extrinsic parameters using the updated focal length.
V-A2 2D Human Pose Estimation
@@
In the implementation, we adopt HRNet [18] and OpenPose [25] to estimate the 2D joint locations of the human and the mirrored human frame by frame from a video sequence of a human moving in front of a mirror. The outputs of HRNet consist of 2D joint locations of eyes, ears, nose, shoulders, elbows, wrists, left and right hips, knees, and ankles. The body25 format output of OpenPose consists of 2D joint locations of eyes, ears, nose, shoulders, elbows, wrists, left and right hips, mid-hip, knees, ankles, and joints on the foot. Since HRNet does not provide 2D joint location of the mid-hip, the hip-angle loss is not adopted when HRNet is utilized. The joints except for those on a face and feet are used to obtain an initial estimation of the camera extrinsic parameters.
V-A3 Training
@@
V-B Baseline Methods and Evaluation Metrics
In order to evaluate the effectiveness of MirrorCalib, we compare its performance with two baseline methods. The first baseline method operates on the same collection of frames and employs the normalized eight-point algorithm without mirror constraints, optimization, or RANSAC. The second baseline method combines the modified eight-point algorithm and optimization with the loss terms in a previous study for multi-view cameras [12] [^1] The method targets uncalibrated and unsynchronized cameras. For a fair comparison, we use synchronized feeds instead.. Note that there is no prior work on calibrating the extrinsic parameters of a virtual camera relative to a real camera using 2D human joint locations.
@@ V-B Baseline Methods and Evaluation Metrics
V-C Datasets
Since there are no existing video datasets with mirrors including ground truth virtual camera parameters, we create a synthetic dataset and collect a real-world dataset.
V-C1 Synthetic Dataset
To create the dataset, we utilize SMPL [19], a differentiable function that produces a human body mesh based on pose and shape parameters. The 3D joint locations can be generated from a linear combination of mesh vertices. We use the SMPL parameters of human motions provided by the SURREAL dataset [27] to obtain the ground-truth 3D body joint locations for the synthetic dataset.
To obtain 2D poses, we project the synthetic 3D body joints onto the image planes of both real and virtual cameras using given intrinsic and extrinsic parameters. These 2D joint locations of the real human and the mirrored human are taken as ground truth. To emulate 2D pose estimation errors, we add Gaussian noise to the ground truth locations and vary its mean and standard deviation. From evaluations using public datasets such as Human3.6 [20] and Mirrored-Human [23], we find the mean and standard deviation errors of state-of-art 2D pose estimators such as HRNet are approximately 0.076 and 4.0 pixels. The resulting noisy 2D joint locations are subsequently used as input to the baseline methods and MirrorCalib. The synthetic dataset allows us to investigate the impact of 2D pose estimation errors and other parameters on different methods in a controlled manner.
@@ V-C1 Synthetic Dataset
V-C2 Real Dataset
[…]
V-D Results on the Synthetic Dataset
V-D1 Impacts of 2D Human Pose Estimation Accuracy
In the first set of experiments, we evaluate the impact of varying errors of 2D human pose estimation on virtual camera extrinsic parameter estimation. In the experiments, the mean error and number of frames (T) are set to be 0.076 pixels and 1000, respectively, but the standard deviation of the errors varies from 1 to 15 pixels. Figure LABEL:extrinsic_experiments shows the rotation and translation errors of extrinsic parameters from different methods. As expected, increasing the standard deviation of 2D human pose estimation leads to larger errors in virtual camera extrinsic parameters in all methods. This underscores the importance of having an accurate 2D human pose estimator.
MirrorCalib outperforms all baseline methods and its performance degrades more gradually as the standard deviation of 2D pose estimation errors increases. When the standard deviation equals 4.0 pixels, the rotation error and translation error of MirrorCalib are 0.62° and 37.3 mm. The inclusion of mirror constraints improves the accuracy of the 8-point algorithm.
V-D2 Impact of the Number of Frames
Next, we assess the impact of T, the number of frames used in virtual camera calibration. The mean and the standard deviation of 2D human pose estimation errors are set to be 0.076 and 4.0 pixels, respectively. As illustrated in Figure LABEL:extrinsic_experiments, increasing the number of frames results in increased accuracy of all methods. However, even when the number of frames is small (e.g., when T = 200), MirrorCalib has achieved a significant reduction in errors. This can be attributed to the ability of the optimization procedure and RANSAC to combat noisy 2D position errors.
V-D3 Ablation Study
We further conduct an ablation study on the synthetic dataset to evaluate the impact of the two pillars of MirrorCalib: optimization and RANSAC. The mean and standard deviation of 2D human pose estimation error are set to be 0.076 and 4.0 pixels respectively and the number of frames equals 1000. The results are summarized in Table I. For comparison, we also include the results of the baseline methods. Table I reveals that the initial camera parameter estimation was significantly impacted by noise. The optimization process can refine the parameters and improve their accuracy. The RANSAC algorithm, as demonstrated in Table I, further increases the accuracy of the result significantly. It can also be observed in the table that the 8-point algorithm with mirror constraints outperforms the general normalized one without mirror constraints.
TABLE I: Ablation study on synthetic dataset. “Init”, “Opt”, and “RANSAC” denote initial, optimization, and RANSAC processes respectively.
Translation Rotation
Error (mm) Error (deg)
Init 393.61 $\pm$ 37.45 6.10 $\pm$2.04
Init + Opt 48.27 $\pm$ 16.31 1.14 $\pm$ 0.83
Init + Opt + RANSAC 37.33 $\pm$ 12.46 0.62 $\pm$ 0.35
Baseline1 482.80 $\pm$ 59.63 7.34$\pm$ 3.52
Baseline2 63.44 $\pm$ 25.81 1.52$\pm$ 0.69
V-D4 Effect of the Accuracy of Intrinsic Parameters
In the previous experiments, we assume the intrinsic parameters are given similar to prior works [11, 12, 13]. When focal length is not known precisely, it is interesting the effect of its accuracy [^2] As we assume the camera has no skew and square pixels and the principal point is located at the center of the image, only focal length is considered in this paper.. Specifically, we set the frame size to 1000, the standard deviation and the mean error of 2D human poses to 4.0 and 0.076 pixels, respectively. We also introduce Gaussian noise to the focal length, assuming a zero mean, and vary the standard deviation of this noise. As shown in Table II and Table III, increasing the error of intrinsic parameters leads to larger extrinsic parameter error. However, the estimation accuracy of extrinsic parameters using MirrorCalib when the focus length error is as high as 100 pixels is still much higher than those in the baseline methods with precise focus length. This shows the robustness of MirrorCalib.
TABLE II: Impact of the accuracy of focal length on rotation estimation (deg).
Methods
Focal Length Error std (pixels)
0 20 40 60 80 100
Baseline1 7.34 $\pm$ 3.52 7.42 $\pm$ 3.71 7.75 $\pm$ 4.18 8.84 $\pm$ 4.03 11.27 $\pm$ 5.22 13.19 $\pm$ 5.34
Baseline2 1.52 $\pm$ 0.69 1.70 $\pm$ 0.84 2.14 $\pm$ 1.07 3.32 $\pm$ 1.21 4.28 $\pm$ 1.67 5.10 $\pm$ 2.13
MirrorCalib 0.62 $\pm$ 0.35 0.71 $\pm$ 0.40 0.98 $\pm$ 0.58 1.61 $\pm$ 0.49 2.53 $\pm$ 0.51 3.46 $\pm$ 0.75
TABLE III: Impact of the accuracy of focal length on translation estimation (mm).
Methods
Focal Length Error std (pixels)
0 20 40 60 80 100
Baseline1 482.80 $\pm$ 59.63 514.77 $\pm$ 54.28 528.92 $\pm$ 60.46 570.17 $\pm$ 72.35 578.82 $\pm$ 63.15 591.48 $\pm$ 77.23
Baseline2 63.44 $\pm$ 25.81 75.89 $\pm$ 27.34 77.53 $\pm$ 28.13 92.84 $\pm$ 34.26 104.31 $\pm$ 42.76 112.56 $\pm$ 49.28
MirrorCalib 37.33 $\pm$ 12.46 49.24 $\pm$ 15.53 62.79 $\pm$ 16.80 70.53 $\pm$ 19.51 78.59 $\pm$ 17.33 94.73 $\pm$ 25.61
V-E Results on the Real Dataset
V-E1 With ground-truth intrinsic parameters
@@
Next, we proceed to conduct experiments on the real dataset to evaluate the effectiveness of MirrorCalib. In this section, we assume the intrinsic parameters are given. The number of frames T is set to be 1000. Table IV shows the results using 2D joint locations detected by HRNet and OpenPose respectively. For both 2D pose estimators, MirrorCalib outperforms baseline methods. Since HRNet is the most state-of-art method compared to OpenPose, using HRNet as our 2D pose estimator results in higher accuracy in the extrinsic parameter estimation.
@@
TABLE IV: Evaluation of virtual camera pose estimation on real dataset using 2D pose locations detected by HRNet and OpenPose.
@@
V-E2 Unknown intrinsic parameters
In this scenario, we follow the steps in Section V-A1 to estimate focus length. The initial error after Step 1) is around 38.8 pixels. By utilizing inliers selected through the RANSAC algorithm after Step 5), a notable improvement can be observed, resulting in a reduced error of 33.9 pixels. As shown in Table V, the utilization of intrinsic parameters estimated from inliers selected by the RANSAC algorithm results in a reduction in extrinsic parameter estimation errors.
TABLE V: Evaluation of virtual camera pose estimation on real dataset using intrinsic parameters estimated from vanishing points.
Translation Rotation
Error (mm) Error (deg)
w/o RANSAC Baseline1 304.19 $\pm$ 52.33 9.21 $\pm$ 2.82
Baseline2 91.63 $\pm$ 28.27 2.63 $\pm$ 1.20
MirrorCalib 73.48 $\pm$ 23.92 2.18 $\pm$ 0.64
with RANSAC Baseline1 294.45 $\pm$ 50.18 9.03 $\pm$ 2.75
Baseline2 86.39 $\pm$ 25.43 2.44 $\pm$ 1.16
MirrorCalib 70.25 $\pm$ 21.36 1.97 $\pm$ 0.49
V-F A Case Study on 3D Human Pose Estimation
One important application of a catadioptric stereo system is 3D human pose estimation. In the literature, many methods exist that given initial 2D poses from multi-view images, estimate 3D poses in an iterative manner [12]. As a case study, we consider a simple pipeline that assumes the intrinsic parameters are known, takes noisy 2D pose estimations as inputs, and calculates 3D joint positions using triangulation using the extrinsic parameters estimated by different methods (Figure 1). The ground-truth 3D joint locations and ground-truth 2D joint locations for this case study are from the synthetic dataset. Gaussian noises are then added to the ground-truth 2D joint locations to obtain noisy 2D joint locations. The mean and standard deviation of noises are set to be 0.076 and 4 pixels. A triangulation algorithm is applied to ground-truth 2D joint locations to obtain the corresponding 3D pose locations using the camera extrinsic parameters from the initial estimation process, the parameters obtained after optimization and RANSAC, and two baseline methods. The performance metric used is PA-MPJPE, which measures the distance (mm) between the predicted 3D key points and their ground truth 3D locations after Procrustes alignment.
As shown in Table VI, 3D pose estimation is significantly more accurate when using the camera extrinsic parameters estimated by MirrorCalib compared to baselines. The result is comparable to SOTA methods for 3D human pose estimation. This experiment also highlights the importance of acquiring precise camera extrinsic parameters for achieving accurate 3D pose estimation.
TABLE VI: Evaluation of 3D human pose estimation using estimated virtual camera extrinsic parameters.

### v3

@@ 3.3.1 Initial Estimation
3.3.2 Refining 2D Joint Locations
@@
Let $\mathbf{u_{i}^{t}}$ and $\mathbf{{u_{i}}^{t}}^{\prime}$ denote the 2D coordinates of joint $i$ and reflected joint $i^{\prime}$ at time $t$, $\mathbf{\Theta}$ and $\mathbf{\Theta^{\prime}}$ denote the extrinsic parameters of camera $C$ and the virtual camera $C^{\prime}$ respectively where $\mathbf{\Theta}$ is known. $\mathbf{\Theta^{\prime}}$ is updated by the modified eight-point algorithm in each iteration, and $\mathbf{X}_{i}^{t}$ denote the 3D coordinates of joint $i$ at time $t$. $\mathbf{X}_{i}^{t}$ can be calculated through triangulation from $\mathbf{u_{i}^{t}}$, $\mathbf{{u_{i}}^{t}}^{\prime}$, $\mathbf{\Theta}$ and $\mathbf{\Theta^{\prime}}$. Our goal is to “denoise” $\mathbf{u}_{i}^{t}$ and ${\mathbf{u}_{i}}^{t^{\prime}}$, to refine the estimation of $\mathbf{\Theta^{\prime}}$.
One body prior is that the length of each limb remains constant over the duration of the recording. Let $l_{k}^{t}$ be the length of the $k_{th}$ bone in the bone set $P$ = {Right Femur, Left Femur, Left Humerus, Right Humerus, Left Ulna, Right Ulna, Left Tibia, Right Tibia, Left Scapular, Right Scaptular, Left Hip, Right Hip} at time $t$ from estimated 2D joint positions. To account for the discrepancies in bone length, we introduce a penalty term that considers the ratio of bone length variation over time to the mean bone length:
@@
Let $\mathbf{v_{ml}}^{t}$ ($\mathbf{v_{mr}}^{t}$) be the vectors from the left (right) hip to the mid-hip. It is expected that the left hip, right hip, and mid-hip joints are co-linear. Thus, we define:
@@
3.3.3 Outlier Rejection
@@
After the optimization in Section 3.3.2, “denoised” 2D joint locations in T frames are obtained. There are many more pairs than needed to solve for the fundamental matrix. To further improve the calibration accuracy, we reject outliers by computing the epipolar distances for eligible pairs for a given fundamental matrix $F$. Let $l(x_{i},C^{\prime})=\mathbf{F}x_{i}$ and $l(x_{i}^{\prime},C)=\mathbf{F}^{\top}{x_{i}}^{\prime}$ be the epipolar lines generated by the 2D point $x_{i}$ on the virtual view $C^{\prime}$ and the matched 2D point ${x_{i}}^{\prime}$ on the real view $C$ respectively [Zhang2014]. Then we can compute the sum of the distance from ${x_{i}}^{\prime}$ to $l(x_{i},C^{\prime})$ and the distance from $x_{i}$ to $l({x_{i}}^{\prime},C)$ in each frame in a video sequence and denote this sum as $g(\mathbf{F})=distance(l(x_{i},C^{\prime}),x_{i}^{\prime})+distance(l(x_{i}^{\prime},C),x_{i})$.
@@
Next, the Random Sample Consensus (RANSAC) algorithm for outlier removal is applied. In each iteration, six pairs of 2D joint positions are selected randomly to calculate the fundamental matrix. This matrix then categorizes the pairs into two groups: inliers and outliers, based on their respective epipolar distances. Specifically, the pairs with an epipolar distance larger than a threshold are defined as outliers. The largest inlier set is used to determine the final reflective fundamental matrix as described in Section 3.1.
4 Evaluation
@@
In this section, we evaluate the performance of MirrorCalib on the collected real-world datasets. Additionally, we conduct a case study on 3D human pose estimation to illustrate the practical application of MirrorCalib.
4.1 Implementation Details
4.1.1 Intrinsic Parameters
@@
MirrorCalib assumes that the intrinsic parameters that characterize the focal length and optical center are given as in previous work [11, 12, 13]. These intrinsic parameters can be obtained from the metadata of videos recorded by professional cameras. Alternatively, they can be estimated using vanishing points [23] under the assumption that the camera has zero skew, square pixels, and the principal point is located at the image center.
4.1.2 2D Human Pose Estimation
@@
In the implementation, we adopt HRNet [18] and OpenPose [25] to estimate the 2D joint locations of the human and the mirrored human frame by frame from a video sequence of a human moving in front of a mirror. Since HRNet does not provide 2D joint location of the mid-hip, the hip-angle loss is not adopted when HRNet is utilized. The joints except for those on the face and feet are used to obtain an initial estimation of the camera extrinsic parameters.
4.1.3 Training
@@
4.2 Baseline Methods and Evaluation Metrics
In order to evaluate the effectiveness of MirrorCalib, we compare its performance with two baseline methods. The first baseline method operates on the same collection of frames and employs the normalized eight-point algorithm [17] without mirror constraints, optimization, or RANSAC. The second baseline method combines the modified eight-point algorithm in initial estimation and optimization using the same strategy in a previous study for multi-view cameras [12] [^1] The method targets uncalibrated and unsynchronized cameras. For a fair comparison, we use synchronized feeds instead.. Note that there is no prior work on calibrating the extrinsic parameters of a virtual camera relative to a real camera using 2D human joint locations.
@@ 4.2 Baseline Methods and Evaluation Metrics
4.3 Datasets
Since there are no existing video datasets with mirrors including ground truth virtual camera parameters, we collect a real-world dataset specified below.
[…]
@@ 4.3 Datasets
4.4 Results
@@
We conduct experiments on the collected dataset to evaluate the effectiveness of MirrorCalib. In this section, we assume the intrinsic parameters are given. The number of frames T is set to be 1000. Table 1 shows the results using 2D joint locations detected by HRNet and OpenPose respectively. For both 2D pose estimators, MirrorCalib outperforms baseline methods. Since HRNet is the most state-of-art method compared to OpenPose, using HRNet as our 2D pose estimator results in higher accuracy in the extrinsic parameter estimation.
@@
Table 1: Evaluation of virtual camera pose estimation on real dataset using 2D pose locations detected by HRNet and OpenPose.
@@
4.5 A Case Study on 3D Human Pose Estimation
One important application of a catadioptric stereo system is 3D human pose estimation. In the literature, many methods exist that given initial 2D poses from multi-view images, estimate 3D poses in an iterative manner [12]. As a case study, we consider a simple pipeline that assumes the intrinsic parameters are known, takes noisy 2D pose estimations as inputs, and calculates 3D joint positions using triangulation using the extrinsic parameters estimated by different methods. We collect a synthetic dataset with ground-truth 3D locations from SMPL [19] models of SURREAL dataset [27]. To obtain 2D poses, we project the synthetic 3D body joints onto the image planes of both real and virtual cameras using given intrinsic and extrinsic parameters. Gaussian noises are then added to the ground-truth 2D joint locations to obtain noisy 2D joint locations. The mean and standard deviation of noises are set to be 0.076 and 4 pixels. A triangulation algorithm is applied to ground-truth 2D joint locations to obtain the corresponding 3D pose locations using the camera extrinsic parameters from the initial estimation process, the parameters obtained after optimization and RANSAC, and two baseline methods. The performance metric used is PA-MPJPE, which measures the distance (mm) between the predicted 3D key points and their ground truth 3D locations after Procrustes alignment.
As shown in Table 2, 3D pose estimation is significantly more accurate when using the camera extrinsic parameters estimated by MirrorCalib compared to baselines. The result is comparable to SOTA methods for 3D human pose estimation. This experiment also highlights the importance of acquiring precise camera extrinsic parameters for achieving accurate 3D pose estimation.
Table 2: Evaluation of 3D human pose estimation using estimated virtual camera extrinsic parameters.